SolveN=8/7Q + 37 for Q
Is the answer
(A) Q=8/7N-37
(B) Q=7/8N-37
(C) Q=7/8(N-37)
(D) Q=8/7(N-37)

Answer by tommyt3rd(5050)

N=8/7Q + 37
N-37=8/7Q + 37-37

8/7 Q = N-37

7/8 8/7 Q = 7/8 (N-37)
Q= (7/8)(N-37) (C)